what is the the answer of 12-{3+[4-5(6-2+7)+1]} in algebra, thank you, [ls help me.

Answer by Theo(13342)

your original equation is shown as:
12-{3+[4-5(6-2+7)+1]}
after converting all symbols to ( and ) and indicating multiplication with *, your equation appears to be:
12 - (3 + (4 - 5 * (6 - 2 + 7) + 1))
the order of arithmetic operations is as follows:
resolve inner parentheses first, then next inner, then next inner, until you have resolved the expression.
the expression itself is assumed to be within an outer set of parentheses (not shown).
within each set of parentheses work from left to right.
if 2 sets of parentheses are on the same level, resolve the left one first and then the right one.
working from left to right within each set of parentheses:
exponents and roots are done first.
multiplies and divides are done next.
adds and subtracts are done next.
your expression is:
12 - (3 + (4 - 5 * (6 - 2 + 7) + 1))
expression within inner set of parentheses appears to be (6 - 2 + 7) so that is resolved first to get 11.
your expression becomes 12 - (3 + (4 - 5 * 11 + 1))
expression within inner set of parentheses now appears to be (4 - 5 * 11 + 1)
working from left to right within that set of parentheses, the multiplication is done first to get (4 - 55 + 1).
the adds and subtracts within that set of parentheses are done next, working from left to right, to get 4 - 55 = -51 + 1 = -50.
your expression becomes 12 - (3 + (-50)) which becomes 12 - (3 - 50).
your next set of inner parentheses now appears to be (3 - 50).
resolve this to get (-47)
your expression becomes 12 - (-47) which becomes 12 + 47 which becomes 59.
your answer should be 59.
i'll go through the process again without all the verbiage so you can see how the order of operations progresses starting with:
12 - (3 + (4 - 5 * (6 - 2 + 7) + 1))
12 - (3 + (4 - 5 * 11 + 1))
12 - (3 + (4 - 55 + 1))
12 - (3 + (-51 + 1))
12 - (3 + (-50))
12 - (3 - 50)
12 - (-47)
12 + 47
59